Calculate the voltage value of a sine wave at 135° whose VRMS= 1000 VAC. Show all work.
1
Expert's answer
2019-04-11T14:19:33-0400

Usually a common sine wave is written as


"V(t)=V_{RMS}\sqrt{2}\text{sin}(\omega t+\phi),"

where "\omega=2\pi f" - angular frequency,

"V_{RMS}\sqrt{2}=V_{amp}" - amplitude (pike or maximum) value of the voltage,

"\phi" - phase angle.


So in our problem the maximum amplitude voltage value is "1000\sqrt{2}=1414.2\text{ V}".